OpenStudy (anonymous):

what is the slope of the line that pass through the given points? (3,2) (5,12) A. 5 B. -5 C. -1/5 D. 1/5

OpenStudy (anonymous):

First of all, we're gonna use the rise over run formula. Do you know what that is?

OpenStudy (anonymous):

The slope's formula is (y2-y1)/(x2-x1). So, we plug in the points (12-2)/(5-3)

OpenStudy (anonymous):

Then you simplify that and solve. (12-2)/(5-3) = 10/2 = ?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

Yea